What is my final grade average if i have a 64 in my math class and got a zero on the exam that was worth 10%..also could you include the steps you took so i can calculate my other classes too.
Answer by dragonwalker(73)

I assume the 64% represents 90% of the grade.
The simplest way is you want the 64 to represent 90% of the whole, or 90 parts,
and zero to represent the remaining 10 parts.
So multiply 64 by 90/100 = 57.6
and 0 by 10/100 = 0
add them together to find the whole:
57.6 + 0 = 57.6 % is the final mark
Ammended:
In response to your question.
You have to look at what each score represents. So 64% is only representing 90% of the total grade. So it only represents 90/100. So you have to multiply 64 by 90/100 to come up with how may points it gives to the grade.
In the same way the 0 represents 10/100 so you multiply the 0 by 10/100 which of course equals 0.
when you add these together to find out how many points for the whole score i.e. 100/100 then you get 57.6 + 0 = 57.6